# GiBy: A Giant-Step Baby-Step Classifier For Anomaly Detection In Industrial Control Systems

### v1

@@ III-C1 Switchboard
III-C2 Linearized state groups
@@
III-C3 Explainability
The switchboard provides bijective mapping from training dataset to testing dataset, using a unique numerical string to represent nn-actuation state. The [LB, UB] bounds are determined from the linearized groups. The bijective nature implies the existence of an inverse mapping, from the testing set to training set (see Fig. 5). I.e., provided with the nn-actuator(s) and the test sensor value, we can determine its corresponding training [LB, UB] bounds for the giant and baby-step. When the sensor value is out of bound, the explainability includes sensor label/name, actuation state, the time index t, and the bound it breached (LB or UB). For the explainability pseudocode refer to lines 6-9 in Algorithm 2.
III-D An example of operations
@@
For other sensors such as FIT201 (see Fig. 4), it is of interest to find the giant and baby-step training bounds. When water is pumped using P101, the flow rate measured by FIT201 settles in a steady range of (mostly) adjacent values. Training directly on the flow rate, and finding bounds, allows us to detect anomalous flows. The baby-step is also of significance w.r.t. this sensor. We are also interested in the bounds on rate of flow change, w.r.t. its nn-neighbor(s) actuation. Later, in section IV we will see that the baby and giant-step may be combined with an extended detection algorithm, to provide up to four different detection mechanisms.
III-E Giant-Step Baby-Step Anomaly Detector
@@
The giant-step baby-step (GiBy) anomaly detector is based on the switchboard implementation, linearization step, and determination of training bounds detailed in Section III-C. The algorithm is run for each sensor with a nn-actuator(s) dependency. We note that the GiBy algorithm is able to take as input, nn-actuator(s) with different actuation sequences. For example, the first actuator might have , with and actuator with . There is also no limitation on the number of input nn-actuators imposed by the algorithm. Algorithm 1 shows GiBy-core training for one sensor. The giant-step training in lines 30-33 involves determining the switchboard states in lines 3-7, linearized state groups (LSG) in lines 15-23 and the bounds in lines 24-29. The baby-step training in lines 34-38 additionally involves computing the in lines 8-14, before it proceeds with the linearization step and bound determination. Algorithm 2 shows GiBy-core testing for one sensor. The giant-step test in lines 10-13 are used to determine the switchboard state and map the test sensor actuation state to its training [LB, UB] bounds (see lines 1-9). The baby-step in lines 14-19 also carries out similar steps, except it is carried out for the test sensor diff value. Note the definition BoundsCheck in lines 1-9 also provides an explanation when an anomaly is detected. The underlined variables are replaced with its actual value when printed.
IV Extending Detection Capabilities
Extended capabilities are built to provide rare event anomaly detection for a time-series window of sensor readings tending towards the boundaries or around the centre of the probability distribution (see Fig. 8). Stealth attacks on ICS were studied in Urbina et al [24]. The goal of an adversary is to keep detection statistic below the selected threshold. Some detection solutions using a residual-error threshold [13] are vulnerable to stealth attacks. We used stateful detection to catch anomalies based on the probability distribution of state values in a time-window. The extended detector is able to flag a subset of time-series sensor readings clustered around (within) the boundaries or towards the centre of the distribution (see Fig. 8); that are rare events not captured during training. Note that it may be used with giant or baby-step. While it improves stealth attack detection (see Section V), this extended detector may not flag attack record readouts exceptionally close to normal operations; typically, when normal operations range, and its statistics are known to an (insider) attacker.
The anomaly probability is deduced using a proposed empirical method. The empirical method is preferred when a relatively large number of entries are available in the training dataset (this is case for the SWaT used within Section II-C). The examples in TABLE IV and TABLE V are hypothetical, and drives the core idea. The extended detector relies on a probability score deduced from the sensor reading.
IV-1 Empirical anomaly score
Consider the hypothetical example in Fig. 7. The range of sensor readings is shown along the x-axis, and the frequency range of these readings is shown up the y-axis. The distribution of sensor values for a given sensor and a specified actuation state is shown. As an example, we begin to calculate the anomaly probability for the sensor reading value of 2 (shown by the red line). We compute four probabilities w.r.t. the sensor — Equation 2 and 3 are the left and right anomaly probabilities, respectively. The empirical anomaly probability in shown in Equation 5, and finally it’s not-anomaly probability is . Let be the sensor reading of interest (for this example ). It is the third reading from the left. Therefore, sensor index . Let be the frequency at the sensor index. Let denote left and for right. We find
For the example with , we have , = 3/36 = 0.083 and = 31/36 = 0.861. The center of the probability distribution is . The anomaly probability for is computed using Equation 5. The edge cases are shown in Algorithm 3.
IV-2 Sliding window in training and testing
The goal is to detect a series of consecutive sensor reading probabilities whose values cluster around the probability boundaries , and (see example Fig. 8). For training, we multiply a series of consecutive not-anomaly probabilities as part of a sliding window, for sensor readings in a time-series dataset. The min and max values of the product are computed. For testing, the same sliding window multiplication is carried out on the test dataset. For the multiplied test probability , when the condition is violated, an anomaly is raised and the explanation provided. The set of consecutive test sensor values for which anomalies are raised are seen as rare events, not captured during training of normal dataset. It may also be used to detect multiple rare events (if any), over a given time interval. TABLE IV shows a hypothetical training dataset for sensor LIT101 and switchboard status 11. Column labeled 1 - PrAnom, is the empirically determined not-anomaly probability. PrAnom is derived using Equation 5. Column labeled SWProduct is the sliding window (sw) product. The sliding window length () was chosen to be 3, for ease of illustration. We compute
| Index | LIT101 | LIT101Diff | sb | 1-PrAnom | SWProduct |
| 1 | 186.2518 | _ | 11 | _ | _ |
| 2 | 186.4088 | 0.1570 | 11 | 0.8 | _ |
| 3 | 186.4099 | 0.0011 | 11 | 0.2 | _ |
| 4 | 186.6050 | 0.1951 | 11 | 0.6 | 0.096 |
| 5 | 186.6835 | 0.0785 | 11 | 0.6 | 0.072 |
| 6 | 187.1546 | 0.4711 | 11 | 0.2 | 0.072 |
For the example in TABLE IV, the min and max bounds for SWProduct are [0.072, 0.096] for .
| Index | LIT101 | LIT101Diff | sb | 1-PrAnom | SWProduct |
| 1 | 111.2324 | _ | 11 | _ | _ |
| 2 | 111.2824 | 0.05 | 11 | 0.4 | _ |
| 3 | 111.3324 | 0.05 | 11 | 0.4 | _ |
| 4 | 111.3824 | 0.05 | 11 | 0.4 | 0.064 |
The example test dataset in TABLE V shows the test sensor readings and its diff. The empirical not-anomaly probability (1-PrAnom) for LIT101Diff is 0.4, since the diffs are the same. The fourth test index with LIT101Diff has sliding window product, SWProduct = 0.064. The verification check therefore is ? Which in our example (SWProduct = 0.064) returns the False condition, and anomaly is flagged by the detector. For the experiments in Section V, the SWaT normal dataset is used for training through use of its sensor values across its linearized group states (see Fig. 5) — for sliding window sizes 5, 10, 25, 50 and 100. For example, a sliding window of size 100 implies that 100 consecutive sensor readings for the given state at one second resolution were trained, and the SWProduct min, max values were the output of that training. This was tested against the SWaT attack dataset. Different window sizes are used to capture anomalies across different time ranges and provide different detection sensitivity depending on for how long the attack lasted. The extended detection training methods are provided in Algorithm 3. The FindMinMaxProduct function at line 37 is used to find the min and max sliding window product bounds for different window lengths and each switchboard state in LSG (derived from training dataset).
For testing the extended algorithm, a similar sequence of steps is used but with the testing dataset. The sensor values used here are the attack dataset readings for giant-step and its diff for baby-step. To find the test sensor value anomaly probability, a binary search is carried out on the corresponding (this is the values sorted in ascending order on frequency, see lines 1-3 in Algorithm 3) derived from extended training. If equals the sorted sensor list value, then PrLeft, PrRight and PrAnom (lines 6-27) are computed for this . For nearest ; on , compute its and on , compute its . I.e., we include the frequency of the tested value in the probability computation. Now that we have the PrLeft and PrRight probabilities for testval, 1 - PrAnom is computed as seen in line 28. The products are then computed for the sliding window lengths and checked against the min, max extended bounds determined in training. An anomaly flag is raised when it is out of bounds. Note that sorting and static probabilities were computed once, stored during training and reused later.
V Experimental Results and Analysis
The experiments are carried out on a Windows 11 machine, with an Intel(R) Core(TM) i7-1195G7 @ 2.90GHz processor, with one of four cores used in experimentation. The GiBy-core detector (Section III-E) and extended algorithm (Section IV) are used. The SWaT normal dataset is used for training and attack dataset is for testing. Table VIII shows the attacks that are detected by GiBy. Attacks 5, 9, 12, 15 and 18 in TABLE VIII had no physical impact on the system. Hence, it is not displayed. For Attack 3, the 1 mm rate of change is within the range of normal behaviour of the tested state. Hence, this rate of change is not flagged for one of two attack states. The extended algorithm can detect the anomaly when the sliding window (multiplied) probability exceeds the seen threshold in training. Attack 13, 14, and 17 are not detected because the changing MV-304 and MV-303 states yet showed normal DPIT-301 readings in attack dataset. Attack 24 on P-203 is not detected because despite turning on this pump, the pH sensor readings were in normal range seen in training. The same is also true with Attack 29 on P-203.
| Giant-step | 219s | 82s | 16s | 407s | 724s |
| Baby-step | 162s | 65s | 15s | 269s | 511s |
| Giant-step | 148s | _ | 148s | 323s/449919 | 0.00071s |
| Baby-step | 145s | 47s | 192s | 320s/449919 | 0.00071s |
The time taken to train and test the normal and attack SWaT datasets are shown in TABLE VI and TABLE VII, respectively. The total training time per sensor was seen to be 9 to 12 minutes. For testing, the total time to copy the nn-actuator(s) and sensor to a new spreadsheet, and computing diff (if any) in the attack dataset are determined in pre-testing steps. The total pre-test time shown is for a total of 449919 attack dataset records. The test step (includes detecting anomaly and providing explanation) is typically seen to be under 1 millisecond per sensor. The time taken for training and pre-testing are linear to the number of records seen in the normal and attack datasets, respectively.
For ECOD and Deep-SVDD in Mathuros et al. [12], the training time is shorter in comparison with GiBy. This is because they train approximately 24 minutes of data, i.e., records for each of the six process stages of SWaT; as opposed to the entire normal dataset of 495000 records trained by GiBy. ECOD detection time was reported to range from 0.72s to 4.62s, whereas for Deep-SVDD, it was 15.88s to 16.45s. Their Integrated Gradient (IG) XAI used with Deep-SVDD was the fastest with a time of 5.33s. However, the XAI provided there use feature ranking to explain what feature likely caused the anomaly. With GiBy, the explanation is precise and pinpoints to the sensor and nn-actuation state (see line 6 in Algorithm 2).
@@ V Experimental Results and Analysis
V-1 Analysis
@@
The attacks in TABLE VIII — 3, 13, 14, 17, 24, 26, 29 were either undetected or partially detected. This is because the range of sensor values in the attack dataset were also seen in the training/normal dataset. However, they are considered as attacks because the sensor readings changed even though it remained below the detection threshold. For example, in attack 29, P-203 is ON and HCl flowed out of the chemical tank and was mixed with water. However, the pH sensor AIT-203 readings remained at levels seen in normal dataset training, though it resulted in wastage of chemical (HCl in this case). It is possible to detect other edge case attack types such as attack 3 and 6; where the sensor reading is either a fixed value or changes by a constant value but remained below the detection threshold. Such a detection may be carried out by observing the number of times the sensor reading pattern repeats and then assigning probabilities for that event. Also, for any detector training there is an entropy loss; where the training dataset input is compressed into a smaller trained output (for example see Fig. 5). For these reasons, we avoid trying to detect specific attacks and designed GiBy to be a generalized one-class classifier. However, the implementer is free to plug-in their edge-case detection models onto GiBy classifier. Without any plug-in, GiBy in TABLE VIII has 33 Y’s and 11 N’s in the Detected column. This translates to a 75% detection of the attacks in the table. Since attack 4 and attack 29 on P-201 has non-existent training data, a correct representation of detected attacks is 33/42, which is 78.5%. We emphasize that the detection accuracy of 78.5% is specific to attacks in TABLE VIII. Accuracy might be alternatively reported as i.) number of detected attack records in SWaT attack dataset divided by the number of total attack records, or worse as ii.) number of total correctly detected labels (Attack/Normal) in attack dataset divided by total number of records in attack dataset. Hence, the interpretation of what accuracy measures and how it is reported may be inconsistent. Next, we use an example attack on SWaT to show the fallibility of anomaly detectors against subtle sensor reading manipulation.
Undetectable attacks
@@
Despite best efforts, a powerful attacker who is able to compromise the PLC may introduce a small constant drift to the sensor reading to avoid detection. Consider the following attack scenario where an attacker has successfully compromised the PLC. Let an attacker manipulate the level sensor LIT-101 reading with the goal to underflow the tank and dry pump using P-101 (see Fig. 4). Further, let the lower limit for LIT-101 be 10 mm; below which the PLC turns OFF the pump P-101; to prevent dry pumping. For the tank filling state MV101 = OPEN and P101 = OFF, the attacker would spoof the LIT101 value on the PLC to ; at each time step . This causes the water level to appear to be higher than it actually is. For the tank draining state MV101 = CLOSE and P101 = ON, the attacker would spoof the LIT101 value on the PLC to ; again, causing the tank to appear fuller than it is. By adjusting the value of and the number of time-steps , it results in tank under-flow and P-101 dry pumping. For example, consider LIT-101 value is 12 mm. Let and . When tank is in filling state for 150 time-steps, . Our accumulated drift is now 1.5 mm. Next, let the tank be in draining state for the next 50-time steps. Then, . This causes additional drift of 0.5, and a total drift of mm. As the tank continues to drain, and LIT-101 sensor reading arrives at 11.9 mm, in reality the tank level is at 9.9 mm. The PLC control logic considers the tank level to be above the set lower-bound, but tank underflow occurred, and P-101 is dry pumping. To make detection further intractable, the drift may be further reduced and the time-steps increased. Further, this attack may be used on one or more sensors simultaneously and might be used to create synchronized failures when additional system information is available. Such subtle attacks tailored to remain below the detection threshold are difficult to detect (refer to Section VIII for possible remedial steps).
VI Discussion
VI-A Detector limitations

### v3

@@ 3.3.1 Switchboard
3.3.2 Linearized state groups
@@
3.3.3 Explainability
The switchboard provides bijective mapping from training dataset to testing dataset, using a unique numerical string to represent nn-actuation state. The [LB, UB] bounds are determined from the linearized groups. The bijective nature implies the existence of an inverse mapping, from the testing set to training set (see Fig. 6). I.e., provided with the nn-actuator(s) and the test sensor value, we can determine its corresponding training [LB, UB] bounds for the giant and baby-step. When the sensor value is out of bounds, the explainability includes sensor label/name, actuation state, the time index t, and the bound it breached (LB or UB). Pseudo-code for the explainability method is included within Algorithm 2.
3.4 An example of operations
@@
For other sensors such as FIT201 (see Fig. 5), it is of interest to find the giant and baby-step training bounds. When water is pumped using P101, the flow rate measured by FIT201 settles in a steady range of (mostly) adjacent values. Training directly on the flow rate, and finding bounds, allows us to detect anomalous flows. The baby-step is also of significance w.r.t. this sensor. We are also interested in the bounds on rate of flow change, w.r.t. its nn-neighbor(s) actuation. Later, in section 4 we will see that the baby and giant-step may be combined with an extended detection algorithm, to provide up to four different detection mechanisms.
3.5 Giant-Step Baby-Step Anomaly Detector
@@
The Giant-step Baby-step (GiBy) anomaly detector is based on the switchboard implementation, linearization step, and determination of training bounds detailed in Section 3.3. The algorithm is run for each sensor with a nn-actuator(s) dependency. We note that the GiBy algorithm is able to take as input, nn-actuator(s) with different actuation sequences. For example, the first actuator might have , with and actuator with . There is also no limitation on the number of input nn-actuators imposed by the algorithm. Both, Algorithm 1 and Algorithm 2 shows GiBy-core training and testing, respectively, for one sensor. Using an example, the giant-step training is visually explained in Fig. 8, the baby-step training in Fig. 9, and the testing and explainability in Fig. 10.
@@
4 Extending Detection Capabilities
Extended capabilities are built to provide anomaly detection for a time-series window of sensor readings tending towards the boundaries or around the centre of the probability distribution (see Fig. 12). Stealth attacks on ICS were studied in Urbina et al [24]. The goal of an adversary is to keep detection statistic below the selected threshold. Some detection solutions using a residual-error threshold [13] are vulnerable to stealth attacks. We used stateful detection to catch anomalies based on the probability distribution of state values in a time-window. The extended detector is able to flag a subset of time-series sensor readings clustered around (within) the boundaries or towards the centre of the distribution (see Fig. 12); those are rare events not captured during training. Note that it may be used with giant or baby-step. While it improves stealth attack detection (see Section 5), this extended detector may not flag attack record readouts exceptionally close to normal operations; typically, this would require an (insider) attacker with extensive prior knowledge of the system inner workings.
The anomaly probability is deduced using a proposed empirical method. The empirical method is preferred when a relatively large number of entries are available in the training dataset (this is case for the SWaT used within Section 2.3). The examples in TABLE V and TABLE VI are hypothetical, and drives the core idea. The extended detector relies on a probability score deduced from the sensor reading.
4.0.1 Empirical anomaly score
Consider the hypothetical example in Fig. 11. The range of sensor readings is shown along the x-axis, and the frequency range of these readings is shown up the y-axis. The distribution of sensor values for a given sensor and a specified actuation state is shown. As an example, we begin to calculate the anomaly probability for the sensor reading value of 2 (shown by the red line). We compute four probabilities w.r.t. the sensor — Equations (2) and (3) are the left and right anomaly probabilities, respectively. The empirical anomaly probability in shown in Equation (5), and finally its not-anomaly probability is . Let be the sensor reading of interest (for this example ). It is the third reading from the left. Therefore, sensor index . Let be the frequency at the sensor index. Let denote left and for right. Equations (1)-(3) are shown below.
For the example with , we have , = 3/36 = 0.083 and = 31/36 = 0.861. The center of the probability distribution is . The anomaly probability for is computed using Equation (5). The edge cases are shown in Algorithm 3.
4.0.2 Sliding window in training and testing
The goal is to detect a series of consecutive sensor reading probabilities whose values cluster around the probability boundaries , and (see example Fig. 12). For training, we multiply a series of consecutive not-anomaly probabilities as part of a sliding window, for sensor readings in a time-series dataset. The min and max values of the product are computed. For testing, the same sliding window multiplication is carried out on the test dataset. For the multiplied test probability , when the condition is violated, an anomaly is raised and the explanation provided. The set of consecutive test sensor values for which anomalies are raised are seen as rare events, not captured during training of normal dataset. TABLE V shows a hypothetical training dataset for sensor LIT101 and switchboard status 11. Column labeled 1 - PrAnom, is the empirically determined not-anomaly probability. PrAnom is derived using Equation (5). Column labeled SWProduct is the sliding window (sw) product. The sliding window length () was chosen to be 3, for ease of illustration. We compute Equation (4).
| Index | LIT101 | LIT101Diff | sb | 1-PrAnom | SWProduct |
| 1 | 186.2518 | _ | 11 | _ | _ |
| 2 | 186.4088 | 0.1570 | 11 | 0.8 | _ |
| 3 | 186.4099 | 0.0011 | 11 | 0.2 | _ |
| 4 | 186.6050 | 0.1951 | 11 | 0.6 | 0.096 |
| 5 | 186.6835 | 0.0785 | 11 | 0.6 | 0.072 |
| 6 | 187.1546 | 0.4711 | 11 | 0.2 | 0.072 |
For the example in TABLE V, the min and max bounds for SWProduct are [0.072, 0.096] for .
| Index | LIT101 | LIT101Diff | sb | 1-PrAnom | SWProduct |
| 1 | 111.2324 | _ | 11 | _ | _ |
| 2 | 111.2824 | 0.05 | 11 | 0.4 | _ |
| 3 | 111.3324 | 0.05 | 11 | 0.4 | _ |
| 4 | 111.3824 | 0.05 | 11 | 0.4 | 0.064 |
| Giant-step | 219s | 82s | 16s | 407s | 724s |
| Baby-step | 162s | 65s | 15s | 269s | 511s |
| Giant-step | 148s | _ | 148s | 323s/449919 | 0.00071s |
| Baby-step | 145s | 47s | 192s | 320s/449919 | 0.00071s |
The example test dataset in TABLE VI shows the test sensor readings and its diff. The empirical not-anomaly probability (1-PrAnom) for LIT101Diff is 0.4, since the diffs are the same. The fourth test index with LIT101Diff has sliding window product, SWProduct = 0.064. The verification check therefore is ? Which in our example (SWProduct = 0.064) returns the False condition, and anomaly is flagged by the detector. For the experiments in Section 5, the SWaT normal dataset is used for training through use of its sensor values across its linearized group states (see Fig. 6) — for sliding window sizes 5, 10, 25, 50 and 100. For example, a sliding window of size 100 implies that 100 consecutive sensor readings for the given state at one second resolution were trained, and the SWProduct min, max values were the output of that training. This was tested against the SWaT attack dataset. Different window sizes are used to capture anomalies across different time ranges and provide different detection sensitivity depending on for how long the attack lasted. The extended detection training methods are provided in Algorithm 3. The FindMinMaxProduct function at line 49 is used to find the min and max sliding window product bounds for different window lengths and each switchboard state in LSG (derived from training dataset).
For testing the extended algorithm, a similar sequence of steps is used but with the testing dataset. The sensor values used here are the attack dataset readings for giant-step and its diff for baby-step. To find the test sensor value anomaly probability, a binary search is carried out on the corresponding (this is the values sorted in ascending order on frequency, in Algorithm 3) derived from extended training. If equals the sorted sensor list value, then PrLeft, PrRight and PrAnom are computed for this . For nearest ; on , compute its and on , compute its . I.e., we include the frequency of the tested value in the probability computation. Now that we have the PrLeft and PrRight probabilities for testval, 1 - PrAnom is computed. The products are then computed for the sliding window lengths and checked against the min, max extended bounds determined in training. An anomaly flag is raised when it is out of bounds. Note that sorting and static probabilities were computed once, stored during training and reused later.
5 Experimental Results and Analysis
The experiments are carried out on a Windows 11 machine, with an Intel(R) Core(TM) i7-1195G7 @ 2.90GHz processor, single core (of 4-total) was used for all experimentation presented. The GiBy-core detector (Section 3) and extended algorithm (Section 4) are used. The SWaT normal dataset is used for training and attack dataset is for testing. TABLE IV shows the attacks that are detected by GiBy. Attacks 5, 9, 12, 15 and 18 in TABLE IV had no physical impact on the system. Hence, they are not displayed. For Attack 3, the 1 mm rate of change is within the range of normal behaviour of the tested state. Hence, this rate of change is not flagged for one of two attack states. The extended algorithm can detect the anomaly when the sliding window (multiplied) probability exceeds the seen threshold in training. Attack 13, 14, and 17 are not detected because the changing MV-304 and MV-303 states all showed normal DPIT-301 readings in attack dataset. Attack 24 on P-203 is not detected because despite turning on this pump, the pH sensor readings were in the normal range seen in training. The same is also true with Attack 29 on P-203.
The time taken to train and test the normal and attack SWaT datasets are shown in TABLE VII and TABLE VIII, respectively. The total training time per sensor was seen to be 9 to 12 minutes. For testing, the total time to copy the nn-actuator(s) and sensor to a new spreadsheet, and computing diff (if any) in the attack dataset are determined in pre-testing steps. The total pre-test time shown is for a total of 449919 attack dataset records. The test step (includes detecting anomaly and providing explanation) is typically seen to be under 1 millisecond per sensor. The time taken for training and pre-testing are linear to the number of records seen in the normal and attack datasets, respectively.
For ECOD and Deep-SVDD in Mathuros et al. [12], the training time is shorter in comparison with GiBy. This is because they train approximately 24 minutes of data, i.e., records for each of the six process stages of SWaT; as opposed to the entire normal dataset of 495000 records trained by GiBy. ECOD detection time was reported to range from 0.72s to 4.62s, whereas for Deep-SVDD, it was 15.88s to 16.45s. Their Integrated Gradient (IG) XAI used with Deep-SVDD was the fastest with a time of 5.33s. However, the XAI provided there ranks the features in non-increasing order of likelihood; to explain what feature most likely caused the anomaly. With GiBy, the explanation is precise and pinpoints to the sensor and nn-actuation state (see line 6 in Algorithm 2).
5.0.1 Analysis
@@
The attacks in TABLE IV — 3, 13, 14, 17, 24, 26, 29 were either undetected or partially detected. This is because the range of sensor values in the attack dataset were also seen in the training/normal dataset. However, they are considered as attacks because the sensor readings changed even though it remained below the detection threshold. For example, in attack 29, P-203 is ON and HCl flowed out of the chemical tank and was mixed with water. However, the pH sensor AIT-203 readings remained at levels seen in normal dataset training, though it resulted in wastage of chemical (HCl in this case). It is possible to detect other edge case attack types such as attack 3 and 6; where the sensor reading is either a fixed value or changes by a constant value but remained below the detection threshold. Such a detection may be carried out by observing the number of times the sensor reading pattern repeats and then assigning probabilities for that event. Also, for any detector training there is an entropy loss; where the training dataset input is compressed into a smaller trained output (for example see Fig. 6). For these reasons, we avoid trying to detect specific attacks and designed GiBy to be a generalized one-class classifier. However, the implementer is free to plug-in their edge-case detection models onto GiBy classifier. Without any plug-in, GiBy in TABLE IV has 33 Y’s and 11 N’s in the Detected column. This translates to a 75% detection of the attacks in the table. Since attack 4 and attack 29 on P-201 has non-existent training data, a correct representation of detected attacks is 33/42, which is 78.5%. We emphasize that the detection accuracy of 78.5% is specific to attacks in TABLE IV. An adjusted detection accuracy presented in the abstract is explained and justified in Section 6.6. Next, we use an example attack on SWaT to show the fallibility of all anomaly detectors against subtle sensor reading manipulation.
5.0.2 Undetectable attacks
@@
Despite best efforts, a powerful attacker who is able to compromise the PLC may introduce a small constant drift to the sensor reading to avoid detection. Consider the following attack scenario where an attacker has successfully compromised the PLC. Let an attacker manipulate the level sensor LIT-101 reading with the goal to underflow the tank and dry pump using P-101 (see Fig. 5). Further, let the lower limit for LIT-101 be 10 mm; below which the PLC turns OFF the pump P-101; to prevent dry pumping. For the tank filling state MV101 = OPEN and P101 = OFF, the attacker would spoof the LIT101 value on the PLC to ; at each time step . This causes the water level to appear to be higher than it actually is. For the tank draining state MV101 = CLOSE and P101 = ON, the attacker would spoof the LIT101 value on the PLC to ; again, causing the tank to appear fuller than it is. By adjusting the value of and the number of time-steps , it results in tank under-flow and P-101 dry pumping. For example, consider LIT-101 value is 12 mm. Let and . When tank is in filling state for 150 time-steps, . Our accumulated drift is now 1.5 mm. Next, let the tank be in draining state for the next 50-time steps. Then, . This causes additional drift of 0.5, and a total drift of mm. As the tank continues to drain, and LIT-101 sensor reading arrives at 11.9 mm, in reality the tank level is at 9.9 mm. The PLC control logic considers the tank level to be above the set lower-bound, but tank underflow occurred, and P-101 is dry pumping. To make detection further intractable, the drift may be further reduced and the time-steps increased. Further, this attack may be propagated on one or more sensors simultaneously and might be used to create synchronized failures when additional system information is available. Such subtle attacks tailored to remain below the detection threshold are difficult to detect (refer to Section 8 for possible remedial steps).
6 Discussion
6.1 Detector limitations
Our core detection solution (see Section 3) is limited to its nn-actuators, for any given sensor. This is not necessarily true when the sensor-actuators(s) relationship are available from the system design document, or when a system expert is able to deduce those relationships. In our work, we assumed such additional information is unavailable.